OpenStudy (anonymous):

the sum of the two smallest of five consecutive odd integers is 50 less than the sum of the other three integers. What is the smallest integer?

OpenStudy (mertsj):

The five integers are: x x+2 x+4 x+6 x+8

OpenStudy (mertsj):

The sum of the two smallest is x + x+2 or 2x+2

OpenStudy (mertsj):

That sum, 2x+2 is 50 less than the sum of the other three. The sum of the other three is 3x+18 So we have 2x+2=3x+18-50

OpenStudy (mertsj):

Can you solve that?

OpenStudy (anonymous):

ya give me a min and I'll tell you the answer

OpenStudy (anonymous):

x=-34

OpenStudy (mertsj):

I got 34 but that is not an odd integer.

OpenStudy (mertsj):

Are you sure the problem said odd?

OpenStudy (anonymous):

my mistake it says even

OpenStudy (mertsj):

Did you find you mistake in getting -34?

OpenStudy (anonymous):

3x+18-50=2x+2

OpenStudy (anonymous):

3x+34=2x

OpenStudy (anonymous):

34=-1x

OpenStudy (anonymous):

x=-34

OpenStudy (mertsj):

3x+18-50=2x+2 3x-32 = 2x=2 3x=2x+34 x=34

OpenStudy (anonymous):

you didn't move the two from the other side cuz 18-50 is -32 then plus 2 is -30

OpenStudy (anonymous):

wait I see what you did

OpenStudy (mertsj):